OpenStudy (anonymous):

The Acceleration function (in m/s^2) and initial velocity for a particle moving along a line is given by a(t)= -14t-21, v(0)=28, 0

OpenStudy (irishboy123):

firstly, checking what you have done already: \[a(t)= -14t-21\] \[v(t)= -7t^2-21t + A\] \[v(0) = 28 = A\] \[v(t)= -7t^2-21t + 28\] next, in terms of distance travelled, you integrate again: \[x(t) = -\frac{7}{3}t^3-\frac{21}{2}t^2 + 28t + B\] and in the first instance the answer might be x(3) - x(0) = ??? the B's cancel, nice. BUT that will give you the difference between the positions or **displacements** of x at t=3 and t=0, not necessarily the distance it has travelled in getting from one displacement to the other --- as the particle could change direction of travel. note it is accelerating in the neg x direction during the whole period. you need to look at the position where v(t) = 0, that is => \[t^2 + 3t - 4 = 0 ; \space{ } t = -4, 1\] and then you need to add together the distances it travels in opposite directions during period 0≤t≤1 and 1≤t≤3, to get the total distance it has travelled.